# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ C-B1 The evaluation of $U_{1}$
By substituting (C-B1) into (113) to replace $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$, we can solve $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ to obtain
@@
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=(\frac{M\delta_{z}\omega_{z}\overline{\omega}_{z}-M\delta_{z}\omega_{z}^{2}\overline{\omega}_{z}^{2}}{1+\delta_{z}\overline{\omega}_{z}^{2}})\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{1-\frac{M\omega_{z}\overline{\omega}_{z}}{L}}{1+\delta_{z}\overline{\omega}_{z}^{2}}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (114)
@@
$-\frac{\jmath u\sqrt{n}}{\sqrt{M^{3}}}\frac{\delta_{z}\overline{\omega}_{z}}{1+\delta_{z}\overline{\omega}_{z}^{2}}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\sqrt{n}}{\sqrt{M}L(1+\delta_{z}\overline{\omega}_{z}^{2})}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u))).$
@@
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ $=\frac{1}{z+\frac{M\omega_{z}\overline{\omega}_{z}^{2}}{L\delta_{z}(1+\delta_{z}\overline{\omega}_{z}^{2})}}\mathbb{E}[(N-M\omega_{z}^{2}\overline{\omega}_{z}^{2}-\frac{M\delta_{z}\omega_{z}\overline{\omega}_{z}^{4}}{1+\delta_{z}\overline{\omega}_{z}^{2}})-\frac{\jmath u\overline{\omega}_{z}}{M(1+\delta_{z}\overline{\omega}_{z}^{2})}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}$ (115)
$-\frac{\jmath u\overline{\omega}_{z}^{2}}{L(1+\delta_{z}\overline{\omega}_{z}^{2})}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,z))$
@@
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=L\delta_{z}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u}{z+\frac{M\omega_{z}\overline{\omega}_{z}^{2}}{L\delta_{z}(1+\delta_{z}\overline{\omega}_{z}^{2})}}[-\frac{\overline{\omega}_{z}}{M(1+\delta_{z}\overline{\omega}_{z}^{2})}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}$ (116)
$-\frac{\overline{\omega}_{z}^{2}}{L(1+\delta_{z}\overline{\omega}_{z}^{2})}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}]\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM})$
$=[M\overline{C}(z)^{\prime}+\frac{N}{z}]\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\sqrt{n}}{\sqrt{M}}[-\frac{\delta\overline{\omega}\overline{\omega}_{z}^{\prime}}{1+\delta\overline{\omega}\overline{\omega}_{z}}-\frac{M\overline{\omega}\omega_{z}\overline{\omega}_{z}^{\prime}}{L\Delta_{\sigma^{2}}(z)\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}}$
$-\frac{M\overline{\omega}\overline{\omega}_{z}}{\Delta_{\sigma^{2}}(z)L(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{[\delta_{z}-(\delta_{z}+\delta_{z}^{2}\overline{\omega}_{z}^{2})]\omega_{z}^{\prime}}{\delta_{z}^{2}}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM}).$
$=[M\overline{C}(z)^{\prime}+\frac{N}{z}]\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\sqrt{n}}{\sqrt{M}}[-\log(1+\delta\overline{\omega}\overline{\omega}_{z})^{\prime}-\frac{M\overline{\omega}\omega_{z}}{\Delta_{\sigma^{2}}(z)L\delta_{z}}\left(\frac{\overline{\omega}_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}\right)^{\prime}$
$-\frac{M\overline{\omega}\overline{\omega}_{z}}{\Delta_{\sigma^{2}}(z)L(1+\delta\overline{\omega}\overline{\omega}_{z})}(\frac{\omega_{z}}{\delta_{z}})^{\prime}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM})$
$=[M\overline{C}(z)^{\prime}+\frac{N}{z}]\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\sqrt{n}}{\sqrt{M}}[-\log(1+\delta\overline{\omega}\overline{\omega}_{z})^{\prime}-[\log(\Delta_{\sigma^{2}}(z))]^{\prime}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM}).$
@@
$U_{1,1}$ $=\jmath\sqrt{\frac{n}{M}}\int_{\sigma^{2}}^{\infty}\frac{N\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{z}-\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\mathrm{d}z$ (118)
$=\jmath\sqrt{nM}\overline{C}(\sigma^{2})\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{nu}{M}[\log(1+\delta\overline{\omega}\overline{\omega}_{z})|_{\sigma^{2}}^{\infty}+\log(\Delta_{\sigma^{2}}(z))|_{\sigma^{2}}^{\infty}]\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$=\jmath\sqrt{nM}\overline{C}(\sigma^{2})\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}-\frac{un}{M}[-\log(\frac{1+\delta\overline{\omega}^{2}}{1+\delta\overline{\omega}})-\log(\Delta_{\sigma^{2}})+\log(\frac{N}{L\delta})]\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,\sigma^{2})+\frac{\mathcal{P}(\frac{1}{\sigma^{2}})u}{\sigma^{2}M})$
$=\jmath\sqrt{nM}\overline{C}(\sigma^{2})\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}-\frac{un}{M}(-\log(\Xi))\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(f({\mathbb{A}},u,\sigma^{2})+\frac{u}{M}).$
@@
$\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\sum_{i,j}\mathbb{E}[{\mathbb{Y}}]_{j,i}^{*}[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\mathbb{E}[\frac{1}{M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}\operatorname{Tr}{\mathbb{A}}-\frac{1}{M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}$ (119)
$+\frac{\jmath u\sqrt{n}}{M^{\frac{3}{2}}}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}-\frac{\jmath un}{\sqrt{Mn}M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}+\frac{\jmath un}{\sqrt{Mn}}\frac{\operatorname{Tr}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}{\mathbb{Q}}}{M}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$-\mathbb{E}\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{2M}+\mathbb{E}\frac{\jmath u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{3M}.$
@@
$\mathrm{Var}(\frac{1}{M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{3}})={\mathcal{O}}(\frac{1}{M}),$ (120)
@@
$\mathrm{Var}(\frac{1}{M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})={\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{3}})={\mathcal{O}}(\frac{1}{M}),$ (121)
@@
$\frac{1}{M}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}=\frac{1}{M^{2}}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}\operatorname{Tr}{\mathbb{A}}-\frac{1}{M^{2}}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}$ (122)
$=-\frac{\mathrm{Cov}(\frac{1}{M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H},\frac{1}{M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{1+\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}}={\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{5}{2}}})={\mathcal{O}}(\frac{1}{M^{\frac{3}{2}}}).$
@@
$\frac{1}{M}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}={\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{5}{2}}}),$ (123)

### 09 Apr 2024 v5

@@ C-B1 The evaluation of $U_{1}$
where $f({\mathbb{A}},u,z)=f({\mathbb{A}},u)\frac{\mathcal{P}(\frac{1}{z})}{z}$. By substituting (C-B1) into (113) to replace $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$, we can solve $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ to obtain
@@
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\Bigl{(}\frac{M\delta_{z}\omega_{z}\overline{\omega}_{z}-M\delta_{z}\omega_{z}^{2}\overline{\omega}_{z}^{2}}{1+\delta_{z}\overline{\omega}_{z}^{2}}\Bigr{)}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{1-\frac{M\omega_{z}\overline{\omega}_{z}}{L}}{1+\delta_{z}\overline{\omega}_{z}^{2}}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (114)
@@
$-\frac{\jmath u\sqrt{n}}{\sqrt{M^{3}}}\frac{\delta_{z}\overline{\omega}_{z}}{1+\delta_{z}\overline{\omega}_{z}^{2}}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{\jmath u\sqrt{n}}{\sqrt{M}L(1+\delta_{z}\overline{\omega}_{z}^{2})}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}(f({\mathbb{A}},u))).$
@@
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ $=\frac{1}{z+\frac{M\omega_{z}\overline{\omega}_{z}^{2}}{L\delta_{z}(1+\delta_{z}\overline{\omega}_{z}^{2})}}\mathbb{E}\Bigl{[}[(N-M\omega_{z}^{2}\overline{\omega}_{z}^{2}-\frac{M\delta_{z}\omega_{z}\overline{\omega}_{z}^{4}}{1+\delta_{z}\overline{\omega}_{z}^{2}})-\frac{\jmath u\overline{\omega}_{z}}{M(1+\delta_{z}\overline{\omega}_{z}^{2})}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})$ (115)
$-\frac{\jmath u\overline{\omega}_{z}^{2}}{L(1+\delta_{z}\overline{\omega}_{z}^{2})}\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}})]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\Bigr{]}+{\mathcal{O}}(f({\mathbb{A}},u,z))$
@@
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=L\delta_{z}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{\jmath u}{z+\frac{M\omega_{z}\overline{\omega}_{z}^{2}}{L\delta_{z}(1+\delta_{z}\overline{\omega}_{z}^{2})}}[-\frac{\overline{\omega}_{z}}{M(1+\delta_{z}\overline{\omega}_{z}^{2})}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]$
$-\frac{\overline{\omega}_{z}^{2}}{L(1+\delta_{z}\overline{\omega}_{z}^{2})}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}})]]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM}\Bigr{)}$
$=[M\overline{C}(z)^{\prime}+\frac{N}{z}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{\jmath u\sqrt{n}}{\sqrt{M}}[-\frac{\delta\overline{\omega}\overline{\omega}_{z}^{\prime}}{1+\delta\overline{\omega}\overline{\omega}_{z}}-\frac{M\overline{\omega}\omega_{z}\overline{\omega}_{z}^{\prime}}{L\Delta_{\sigma^{2}}(z)\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}}$
$-\frac{M\overline{\omega}\overline{\omega}_{z}}{\Delta_{\sigma^{2}}(z)L(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{[\delta_{z}-(\delta_{z}+\delta_{z}^{2}\overline{\omega}_{z}^{2})]\omega_{z}^{\prime}}{\delta_{z}^{2}}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}\Bigl{(}f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM}\Bigr{)}.$
$=[M\overline{C}(z)^{\prime}+\frac{N}{z}]\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+\frac{\jmath u\sqrt{n}}{\sqrt{M}}[-\log(1+\delta\overline{\omega}\overline{\omega}_{z})^{\prime}-\frac{M\overline{\omega}\omega_{z}}{\Delta_{\sigma^{2}}(z)L\delta_{z}}\left(\frac{\overline{\omega}_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}\right)^{\prime}$ (116)
$-\frac{M\overline{\omega}\overline{\omega}_{z}}{\Delta_{\sigma^{2}}(z)L(1+\delta\overline{\omega}\overline{\omega}_{z})}(\frac{\omega_{z}}{\delta_{z}})^{\prime}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}\Bigl{(}f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM}\Bigr{)}$
$=[M\overline{C}(z)^{\prime}+\frac{N}{z}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{\jmath u\sqrt{n}}{\sqrt{M}}[-\log(1+\delta\overline{\omega}\overline{\omega}_{z})^{\prime}-[\log(\Delta_{\sigma^{2}}(z))]^{\prime}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}\Bigl{(}f({\mathbb{A}},u,z)+\frac{\mathcal{P}(\frac{1}{z})u}{zM}\Bigr{)}.$
@@
$U_{1,1}$ $=\jmath\sqrt{\frac{n}{M}}\int_{\sigma^{2}}^{\infty}\frac{N\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]}{z}-\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]\mathrm{d}z$ (118)
$=\jmath\sqrt{nM}\overline{C}(\sigma^{2})\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\frac{nu}{M}[\log(1+\delta\overline{\omega}\overline{\omega}_{z})|_{\sigma^{2}}^{\infty}+\log(\Delta_{\sigma^{2}}(z))|_{\sigma^{2}}^{\infty}]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$
$=\jmath\sqrt{nM}\overline{C}(\sigma^{2})\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]-\frac{un}{M}[-\log(\frac{1+\delta\overline{\omega}^{2}}{1+\delta\overline{\omega}})-\log(\Delta_{\sigma^{2}})+\log(\frac{N}{L\delta})]\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}(f({\mathbb{A}},u,\sigma^{2})+\frac{\mathcal{P}(\frac{1}{\sigma^{2}})u}{\sigma^{2}M})$
$=\jmath\sqrt{nM}\overline{C}(\sigma^{2})\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]-\frac{un}{M}(-\log(\Xi))\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}(f({\mathbb{A}},u,\sigma^{2})+\frac{u}{M}).$
@@
$\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\sum_{i,j}\mathbb{E}[[{\mathbb{Y}}]_{j,i}^{*}[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\mathbb{E}\Bigl{[}[\frac{1}{M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})\operatorname{Tr}({\mathbb{A}})-\frac{1}{M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})$ (119)
$+\frac{\jmath u\sqrt{n}}{M^{\frac{3}{2}}}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})-\frac{\jmath un}{\sqrt{Mn}M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})+\frac{\jmath un}{\sqrt{Mn}}\frac{\operatorname{Tr}({\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}{\mathbb{Q}})}{M}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\Bigr{]}$
$-\mathbb{E}\Bigl{[}\frac{u^{2}\partial\nu_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{2M}+\mathbb{E}\frac{\jmath u^{3}\partial\beta_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{\partial Y_{j,i}}\frac{[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}]_{j,i}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}}{3M}\Bigr{]}.$
@@
$\mathrm{Var}(\frac{1}{M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}))={\mathcal{O}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{3}})={\mathcal{O}}(\frac{1}{M}),$ (120)
@@
$\mathrm{Var}(\frac{1}{M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}))={\mathcal{O}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{3}})={\mathcal{O}}(\frac{1}{M}),$ (121)
@@
$\frac{1}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]=\frac{1}{M^{2}}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]\operatorname{Tr}({\mathbb{A}})-\frac{1}{M^{2}}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]$ (122)
$=-\frac{\mathrm{Cov}(\frac{1}{M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}),\frac{1}{M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}))}{1+\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}}={\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{5}{2}}})={\mathcal{O}}(\frac{1}{M^{\frac{3}{2}}}).$
@@
$\frac{1}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})]={\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{5}{2}}}),$ (123)